OpenStudy (anonymous):

Sam bought 2 1/4 pounds of Swiss cheese and 1 1/2pounds of cheddar at the deli. How many pounds of cheese did Sam buy altogether?

OpenStudy (anonymous):

(2 1/4) + (1 1/2) = 3.75

OpenStudy (asnaseer):

correct, although you can express it as:\[2\frac{1}{4}+1\frac{1}{2}=3+\frac{1}{4}+\frac{1}{2}=3+\frac{1}{4}+\frac{2}{4}=3+\frac{3}{4}=3\frac{3}{4}\]